An ecenemist is studying the 1eb market in Denver area neichborhoods. Let x represent the total number aobs in a oven neighborhood, and let y represent the number of ertry-leves jobs in the same neighcorhood. A sampet sx Denwer neschborhoods gave the toilowing intormation (units in huncreds ofobs. a) Draw a seater dagram displaying the dats Flash Payer 0 or higher s required for You can get Rlash Player tree from Adobes website. t) Verty the gven tumsrx, Ζν. D2M2, Ery, and tte van af the tar pte corrlation cetfident r, (Round your vatn trrtoerce do mal plann.) 2y (c Pind x, and men find the aua ond, the least saares ine9-аャa und your answers for x andytotodeanu paces Rund your ans es for a ansutnee dec m려 places. (d) Graph the least-squares line. Be sure to plat the pontas a point on the line 1 20 3400 120 0 405080 20 0 405080 n) Prd tte val n of the e mont of determinat r , what pe antaga c the vanaton in yar, be egan dbyt c r spring variation nrare the ︼ tan sine? what tagn is um aned? Ra dyrarm er r_tthme deci al la es. Rond your answers fr t ต pe certa sto eder s wthx-33 hunered jobs, how many are predictoe to be entry level jobst (Round your answer to bwo decimal places)
Answer #1

S.No 1 2 16 30 50 28 256.0000 900.0000 2500.0000 784.0000 2500.0000 625.0000 EX 7565.0000 9.0000 9.0000 25.0000 25.0000 81.0000 9.0000 48.00 90.00 250.00 140.00 450.00 75.00 25 EX 199.0000 total 28.0000 158.0000 1053.0000 6.0000 33.1667 4.6667 13.8912 2.3381 24.8667 0.7656 s,-(V(Zx-(Zx)2/n)/(n-1)) Covxys(ΣΧΥ_ (ΣΥΣΥ)/n)/(n-1)- r-Cov/(Sx*Sy)- from above slope- 0.129 intercept- o 0.393 hence line equation: 0.393 0.129 x

b)

sumx =199

sumy=28

sumx2=7565

sumy2=158

sumxy=1053

r=0.766

c)

xbar=33.17

ybar=4.67

yhat=0.393+0.129x

d)

top left plot is correct

e)

r2 =0.586

explained =58.6%

unexplained=41.4%

r)

predicted value =0.393+0.129*33=4.65